OpenStudy (anonymous):

if x/y = 3/4 find 5x-3y/7x+2y

OpenStudy (anonymous):

x/y=3/4 or,x=3y/4

OpenStudy (anonymous):

put the value of x at the expression

OpenStudy (anonymous):

3/29